OpenStudy (anonymous):

]What is the slope of the line passing through the point (3, −4) and perpendicular to the line y = 3x + 5? Write your answer as a fraction.

OpenStudy (anonymous):

\[y + 4 = -\frac{1}{3}(x-3)\] lines that are perpendicular have a slope that is the negative reciprocal of each other. So just use point slope formula and solve for y. \[y + 4 = -\frac{1}{3}x + 1\] \[y = -\frac{1}{3}x - 3\]